OpenStudy (anonymous):

What are the x– and y–intercepts of the graph of y = 2x^2 – 8x – 10?

OpenStudy (anonymous):

The axis of symmetry for the first equation would be x = 3. To find the x intercepts of y = 2x^2 - 8x - 10, we need to set the left side to zero and solve: y = 2x^2 - 8x - 10 0 = 2x^2 - 8x - 10 0 = (2x + 2)(x - 5) 2x + 2 = 0 2x = -2 x = -1 x - 5 = 0 x = 5 So, the x intercepts are -1 and 5 To find the y intercepts, we need to set x = 0 in the equation and solve for y: y = 2x^2 - 8x - 10 y = 2(0)^2 - 8(0) - 10 y = -10 So, the y intercept is -10 I leave the last problem to you, since you need to explain why you chose the method that you did. Good luck! MY MEDAL

OpenStudy (anonymous):

thank you so much you are a big help(:

OpenStudy (anonymous):

ANOTHER QUESTION PLEASE CAN HELP YOU

OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = –3(x + 8)^2 + 5?

OpenStudy (anonymous):

It is x=-8 when x=-8+a and x=-8-a y will get the same value = -3a^2 + 5

OpenStudy (anonymous):

thanks